OpenStudy (anonymous):

If f(x)= 3x^2-8 and g(x) = 2x, evaluate f(x) divided by g(x) for x = 2.

OpenStudy (anonymous):

f(x) / g(x) = \[(3x^{2} - 8) / 2x \]

OpenStudy (anonymous):

\[\frac{ 3x }{ 2 } - \frac {4} {x}\] now just plug in x= 2